two torch light batteries each having e.m.f 1.5v and internal resistance of 1 ohms resistor.calculate the current in this resistor if the cells are connected in parallel
1
Expert's answer
2020-12-03T06:58:09-0500

If the cells are in parallel, "e.m.f = 1.5v"

find effective resistance, "R_{E}"

"\frac{1}{R_{E}}=\frac{1}{1}+\frac{1}{1} =2" "\Omega"

"R_E = 1\/2=0.5"

"I = \frac{v}{R_E} = \frac{1.5}{0.5} = 3A"